OpenStudy (princessaurora):

Assuming that the amount of water leaking from the faucet every hour increases uniformly by (1/12) of the amount that leaked in the previous hour, write a function representing the amount of water leaked each hour. *I can't post a pic of the beaker but the fill lien is at 75ml but the beaker goes, 10,20,20,40,50,60,70,80,90,100*

OpenStudy (princeharryyy):

let x denote the hrs water leakage first hr = (1/12)*x at x =1 x varies.

OpenStudy (princeharryyy):

f(x) = (1/12)*x